### v1

# Correlator with tensor currents and two masses at two loops

Abstract
We calculate the vacuum-to-vacuum correlator of two quark tensor currents with two massive quarks, retaining full momentum dependence. For the first time, we include perturbative corrections up to next-to-leading order. Our fully analytical expressions are provided in machine-readable form. Furthermore, we present numerical results for various input parameters, including an estimate of the scale uncertainties. Our results are essential input for applications of dispersive methods, including unitarity bounds and QCD sum rules.
1 Introduction
Flavour physics provides an ideal framework for indirect searches for physics beyond the Standard Model (SM). This is accomplished by comparing precise experimental measurements with the corresponding SM predictions for observables in hadron decays, particularly in the decays of $B$ mesons. In fact, the past three decades have been marked by a monumental effort to reduce both experimental and theoretical uncertainties, thereby enhancing the effectiveness of indirect searches.
The main obstacle to reducing theoretical uncertainties lies in the hadronic matrix elements — such as decay constants and hadronic form factors — whose predictions are governed by low-energy, and thus non-perturbative, QCD dynamics. Techniques based on dispersion relations are extremely useful for predicting or constraining these matrix elements. Indeed, dispersion relations — one of the few exact results known in QCD — relate matrix elements to perturbatively calculable quantities. The two most commonly used techniques based on dispersion relations are unitarity bounds and QCD sum rules; see Refs. [1, 2, 3] for reviews.
Both the derivation of unitarity bounds and QCD sum rules requires the calculation of (at least) a two-point correlator with quark currents. Two-point correlators can be written as:
[…]
where $\psi_{a,b}$ are quark fields and $\Gamma$ denotes a Dirac matrix or a product of Dirac matrices. These correlators are typically calculated using an operator product expansion (OPE), which expresses the correlator as a sum over local operators. Each term consists of a Wilson coefficient, calculable perturbatively in the strong coupling constant $\alpha_{s}$, multiplied by a local operator of definite dimension, capturing non-perturbative effects. This naturally leads to a twofold expansion: in the operator dimensions and in $\alpha_{s}$. In most cases, the dominant contribution to these OPEs comes from the dimension-zero operator, i.e., the identity operator, which requires the calculation of bubble diagrams. For (axial-)vector currents, namely $\Gamma=\gamma^{\mu}(\gamma_{5})$, two-point correlators have been calculated at next-to-leading order (NLO) in $\alpha_{s}$ with two massive quarks and full momentum dependence [4]. These correlators have also been calculated next-to-next-to-leading order (NNLO) with two massive quarks, but only at $q^{2}=0$ using both analytical and numerical methods [5]. For (axial-)tensor currents, namely $\Gamma=\sigma^{\mu\nu}(\gamma_{5})$, two-point correlators have only been calculated at NLO with two massive quarks and $q^{2}=0$ [6]. This correlator has also been computed recently using lattice QCD [7].
In this article, we present the first complete analytical calculation of (axial-)tensor two-point correlators at NLO, with two massive quarks and full momentum dependence. Furthermore, we estimate the scale uncertainties and compare our numerical results with those obtained in the lattice QCD study of Ref. [7].
The remainder of this article is organized as follows. In Section 2, we provide details of our calculation and present the analytical results. In Section 3, we give explicit numerical results for selected values of the quark masses and $q^{2}$. Our conclusions are summarized in Section 4.
2 Analytical calculation
In this section, we first introduce our notation and review the one-loop contribution to the correlator (Subsection 2.1). We then present the two-loop calculation, providing explicit analytical results for the relevant correlators (Subsection 2.2). Finally, we compare our findings with existing results in the literature where they are available (Subsection 2.3).
2.1 Theoretical framework and one-loop result
We consider the correlators
[…]
$=\left(\frac{q^{\mu}q^{\nu}}{q^{2}}-g^{\mu\nu}\right)\Pi_{T}(q^{2})\,$, (2.1)
[…]
$=\left(\frac{q^{\mu}q^{\nu}}{q^{2}}-g^{\mu\nu}\right)\Pi_{AT}(q^{2})\,$, (2.2)
where $\psi_{a,b}$ are massive quark fields and $\sigma^{\mu\alpha}\equiv\frac{i}{2}[\gamma^{\mu},\gamma^{\alpha}]$. Note that the second line in Eqs. (2.1) and (2.2) follows directly from Lorentz covariance. Correlators such as $\Pi_{T}$ and $\Pi_{AT}$ satisfy the following identity, known as the dispersion relation (see, e.g., Refs. [1, 2]):
$\Pi_{T}(q^{2})$ $=\frac{1}{\pi}\int\limits_{0}^{\infty}ds\,\frac{\text{Im}\,\Pi_{T}(s)}{s-q^{2}-i\epsilon}\,.$ (2.3)
These correlators can be calculated using OPE for $q^{2}\ll\left(m_{1}+m_{2}\right)^{2}-\left(m_{1}+m_{2}\right)\Lambda_{\rm QCD}$, with $\Lambda_{\rm QCD}\sim 300\,\text{MeV}$ (see, e.g., Refs. [8, 1]). The leading power contribution to the OPE in the regime defined above can be calculated using standard perturbation theory. In practice, it suffices to compute $\Pi_{T}$, since the results for $\Pi_{AT}$ can be obtained directly by performing the replacement $m_{1}\to-m_{1}$ [4]. For the remainder of this article, we therefore focus exclusively on $\Pi_{T}$, with the understanding that the corresponding results for $\Pi_{AT}$ can be easily derived.
Figure 1: Leading-power contribution to the OPE at LO (a) and NLO (b)-(d). The crossed circles denote the insertion points of the tensor current.
The leading-power contribution to the OPE for $\Pi_{T}$ at leading order in $\alpha_{s}$ can be obtained straightforwardly, as it requires only the calculation of the one-loop diagram shown in Fig. 1(a). Such one-loop calculations are fully automated using, for example, Package X [9] or FeynCalc [10]. This result reads
$\Pi_{T}(q^{2})\big{|}_{\text{1-loop}}$ $=\frac{q^{2}}{8\pi^{2}}(6m_{1}m_{2}+q^{2})\left(\frac{1}{\epsilon}+\log\left(\frac{\mu^{2}}{m_{1}m_{2}}\right)\right)$ (2.4)
$+\frac{1}{48\pi^{2}q^{2}}\Bigg{(}2q^{2}\left(-3q^{2}\left(m_{1}^{2}-12m_{1}m_{2}+m_{2}^{2}\right)-6\left(m_{1}^{2}-m_{2}^{2}\right)^{2}+8(q^{2})^{2}\right)$
$+6\left((m_{1}-m_{2})^{2}-q^{2}\right)\left(2(m_{1}+m_{2})^{2}+q^{2}\right)\lambda^{1/2}\left(m_{1}^{2},m_{2}^{2},q^{2}\right)$
$\times\left(\log(2m_{1}m_{2})-\log\left(\lambda^{1/2}\left(m_{1}^{2},m_{2}^{2},q^{2}\right)+m_{1}^{2}+m_{2}^{2}-q^{2}\right)\right)$
$+6(m_{1}-m_{2})(m_{1}+m_{2})^{3}\left(2(m_{1}-m_{2})^{2}-3q^{2}\right)\log\left(\frac{m_{1}}{m_{2}}\right)\Bigg{)}+\mathcal{O}(\epsilon)\,$,
where $\lambda(m_{1}^{2},m_{2}^{2},q^{2})\equiv\left(m_{1}^{2}-m_{2}^{2}-q^{2}\right)^{2}-4m_{2}^{2}\,q^{2}$ is the Källén function. Note that throughout this work, we perform the substitution
$\mu^{2\epsilon}\to\frac{\mu^{2\epsilon}}{\Gamma(1+\epsilon)}\;,$ (2.5)
which allows us to renormalise the strong coupling in the $\overline{\text{MS}}$ scheme by subtracting only pure poles in $\epsilon$. The explicit result of Eq. (2.4) is useful for illustrating how certain ultraviolet divergences are treated within the framework of dispersive techniques. Typically, instead of renormalizing vacuum-to-vacuum correlators such as $\Pi_{T}$, one subtracts from them the first few terms of their Taylor expansion in $q^{2}$. This is equivalent to taking derivatives of Eq. (2.3) with respect to $q^{2}$:
$\chi_{T}(Q^{2};k)\equiv\frac{1}{k!}\left[\frac{\partial}{\partial q^{2}}\right]^{k}\Pi_{T}(q^{2})\Bigg{|}_{q^{2}=-Q^{2}}$ $=\frac{1}{\pi}\int\limits_{0}^{\infty}\!ds\,\frac{\text{Im}\,\Pi_{T}(s)}{(s+Q^{2})^{k+1}}\,.$ (2.6)
Here, we define the quantity $\chi_{T}$ for later use. It is straightforward to see that in this case at least three derivatives are required, and hence $k\geq 3$. Note that beyond one-loop order, in addition to performing subtractions, the mass, the current, and the strong coupling must still be renormalized to cancel all divergences.
2.2 Two-loop result
We present the first calculation of the NLO corrections to the leading-power contribution in the OPE for $\Pi_{T}$ for two massive quarks and full $q^{2}$ dependence. This involves evaluating the two-loop diagrams shown in Fig. 1(b)-(d).
In this study, the integration-by-parts (IBP) reduction can be performed relatively efficiently with contemporary computational tools, as the problem involves only three independent scales ($m_{1}$, $m_{2}$, and $q^{2}$) and the diagram topologies are comparatively simple, allowing for a manageable number of master integrals. The IBP reduction is carried out using FIRE [11], yielding a total of four master integrals not counting those which can be obtained by swapping the quark flavour labels. All of these master integrals have been evaluated analytically up to the required order in $\epsilon$ in the literature, with explicit expressions available in Ref. [4]. Combining these results yields the unrenormalized expression for $\Pi_{T}$. For completeness we also obtain the expressions for the scalar $\Pi_{S}$ and vector $\Pi_{VL,VT}$ correlators, defined as
[…]
$=\frac{q^{\mu}q^{\nu}}{q^{2}}\Pi_{VL}(q^{2})+\left(\frac{q^{\mu}q^{\nu}}{q^{2}}-g^{\mu\nu}\right)\Pi_{VT}(q^{2})\,.$ (2.8)
We provide these results in machine-readable Mathematica format as ancillary files attached to the arXiv version of this paper. For completeness, we also provide the LO results. These files are named
LO_S.m LO_VL.m LO_VT.m LO_TT.m
NLO_S_unren.m NLO_VL_unren.m NLO_VT_unren.m NLO_TT_unren.m
We now need to renormalize these expressions. At two-loop order, one needs to renormalize the masses and the currents:
$\Pi_{T}=\Pi_{T}\big{|}_{\rm unren.}+\Pi_{T}\big{|}_{\rm CT-mass}+\Pi_{T}\big{|}_{\rm CT-curr}\,$, (2.9)
where the counterterms are defined as
$\Pi_{T}\big{|}_{\rm CT-mass}$ $=-\delta m_{1}\frac{\partial}{\partial m_{1}}\Pi_{T}(q^{2})\big{|}_{\text{1-loop}}-\delta m_{2}\frac{\partial}{\partial m_{2}}\Pi_{T}(q^{2})\big{|}_{\text{1-loop}}\,$, (2.10)
$\Pi_{T}\big{|}_{\rm CT-curr}$ $=2(Z_{T}-1)\Pi_{T}(q^{2})\big{|}_{\text{1-loop}}\,.$ (2.11)
The renormalization constants in the $\overline{\text{MS}}$ scheme at one loop are given by [^1] For the vector and scalar currents the current renormalization constants are given by $Z_{V}=1$ and $Z_{S}=1-3\alpha_{s}/(4\pi\,\epsilon)C_{F}+{\cal O}{\left(\alpha_{s}^{2}\right)}$, respectively.
$\delta m_{i}$ $\equiv m_{i}(1-Z_{m})=m_{i}\frac{3\,\alpha_{s}}{4\,\pi\,\epsilon}C_{F}+{\cal O}{\left(\alpha_{s}^{2}\right)}\,$, (2.12)
$Z_{T}$ $=1+\frac{\alpha_{s}}{4\,\pi\,\epsilon}C_{F}+{\cal O}{\left(\alpha_{s}^{2}\right)}\,$, (2.13)
with $C_{F}=4/3$, while the mass renormalization constant in the on-shell scheme at one loop is given by
$\delta m_{i}\equiv m_{i}(1-Z_{m})$ $=\frac{m_{i}}{\epsilon}\frac{\alpha_{s}}{\pi}\bigg{(}\frac{\mu^{2}}{m_{i}^{2}}\bigg{)}^{\!\!\epsilon}C_{F}\frac{3-2\epsilon}{4-8\epsilon}+{\cal O}{\left(\alpha_{s}^{2}\right)}\,.$ (2.14)
Therefore, using Eqs. (2.9)-(2.14), we obtain the renormalized expressions for the correlators. As with the unrenormalized ones, these are provided in machine-readable Mathematica format as ancillary files attached to the arXiv version of this paper:
NLO_S_MSbar.m NLO_VL_MSbar.m NLO_VT_MSbar.m NLO_TT_MSbar.m
NLO_S_OS.m NLO_VL_OS.m NLO_VT_OS.m NLO_TT_OS.m
For the reader’s convenience, we also provide the explicit expressions for the imaginary parts of these correlators, which can be extracted from the results above:
Im_NLO_S_MSbar.m Im_NLO_VL_MSbar.m Im_NLO_VT_MSbar.m Im_NLO_TT_MSbar.m
Im_NLO_S_OS.m Im_NLO_VL_OS.m Im_NLO_VT_OS.m Im_NLO_TT_OS.m
Working with the imaginary part is often more convenient in the context of unitarity bounds, as taking derivatives — i.e. performing subtractions — then becomes trivial (see Eq. (2.6)). Moreover, in the context of QCD sum rules, knowledge of the imaginary part constitutes a key ingredient [1].
The $\chi_{\Gamma}(Q^{2};k)$ functions (i.e. the moments of the correlator) defined in Eq. (2.6) can in principle be obtained from the results above. However, since the lowest finite moments of the correlators at $Q^{2}=0$ are needed quite commonly, we also provide their explicit expressions in the files222 It can be shown that the lowest admissible values of $k$ are $2$, 11, 22, and $3$ for $\chi_{S}$, $\chi_{VL}$, $\chi_{VT}$, and $\chi_{T}$, respectively.
Moment_LO_S.m Moment_NLO_S_unren.m Moment_NLO_S_MSbar.m Moment_NLO_S_OS.m
Moment_LO_VL.m Moment_NLO_VL_unren.m Moment_NLO_VL_MSbar.m Moment_NLO_VL_OS.m
Moment_LO_VT.m Moment_NLO_VT_unren.m Moment_NLO_VT_MSbar.m Moment_NLO_VT_OS.m
Moment_LO_TT.m Moment_NLO_TT_unren.m Moment_NLO_TT_MSbar.m Moment_NLO_TT_OS.m
2.3 Comparison with previous results
We conclude this section by comparing our results with the literature. As a validation of our calculation, we compare our results for $\Pi_{S}$ and $\Pi_{V}$ with those reported in Refs. [4, 5] and find exact agreement. We compare our results with Ref. [12], which presents the NLO expression for $\text{Im}\,\Pi_{T}$ in the regime $m_{1}\ll m_{2}$, including terms up to ${\cal O}(m_{1}/m_{2})$, and find exact agreement.
Finally, we verify our results against those of Ref. [6], which presents $\Pi_{S}$, $\Pi_{V}$, and $\Pi_{T}$ at $q^{2}=0$ with the minimal number of subtractions (i.e. $\chi_{S}(0;3)$, $\chi_{V}(0;3)$, and $\chi_{T}(0;3)$). While Ref. [6] does not specify the treatment of current renormalization, we find agreement with their results only if the last term in Eq. (2.9) is omitted. When instead including this term, the expression for $\Pi_{V}$ remains unchanged, since the vector current does not require renormalization, whereas our corresponding results for the $\mathcal{O}(\epsilon^{0})$ parts of $\Pi_{S}$ and $\Pi_{T}$ are different and disagree with the results of Ref. [6]. As expected, we find that the moments and the imaginary parts of $\Pi_{S}$ and $\Pi_{T}$ are free of $\epsilon$-poles only when the current renormalization is performed.
Therefore, our result for $\Pi_{T}$ not only supersedes previous literature by providing the full $m_{1}$, $m_{2}$, and $q^{2}$ dependence, but also resolves an inconsistency with earlier results.
3 Numerical results
In this section, we provide explicit numerical results based on the analytical expressions derived in the previous section. Table 1 reports numerical values of the functions $\chi_{T}(Q^{2};k)$ and $\chi_{AT}(Q^{2};k)$ defined in Eq. (2.6). The values at $Q^{2}=0$ and $k=3$ are of particular importance, as they are most commonly used in the context of unitarity bounds [6, 13, 14]. Nevertheless, other values of $Q^{2}$ and $k$ can also be relevant in that context [8, 15].
We adopt the $\overline{\text{MS}}$ scheme for the quark masses, since it provides a short-distance mass definition that avoids the infrared sensitivities inherent in the on-shell scheme. In particular, the pole mass suffers from renormalon ambiguities of order $\Lambda_{\rm QCD}$ making it ill-suited for perturbative analyses of correlators such as those considered in this work [16].333 See also e.g. Ref. [17] for a review of quark-mass renormalisation schemes. In particular, it is argued in Refs. [16, 17] that the kinetic mass scheme is the most appropriate for semi-leptonic decays of heavy-light mesons. However, since we are performing a calculation at a relatively low order in perturbation theory, such considerations likely lie beyond the precision attainable here. In particular, we observe good apparent perturbative convergence through NLO using the $\overline{\text{MS}}$ scheme, which a posteriori justifies this choice. The values of the quark masses used in our numerical analysis are also listed in Table 1. The central values are taken from the Particle Data Group [18], while their uncertainties are neglected, since their effect on the final results is much smaller than that of the scale uncertainties. We then evolve the quark masses to the required scale using the three-loop running implemented in RunDec [19], with $\alpha_{s}(m_{Z})=0.1180$ [18].
Value of the masses $\chi_{\Gamma}(Q^{2};k)$ Central value $\pm$ unc. $\text{GeV}^{-2}$]
$m_{1}=0$
$m_{2}=\overline{m}_{b}(\overline{m}_{b})=4.183\,\text{GeV}$
$\chi_{T}(0;3)=\chi_{AT}(0;3)$
$(4.060_{-0.589}^{+0.119})\cdot 10^{-4}$
$\chi_{T}(Q_{b}^{2};3)=\chi_{AT}(Q_{b}^{2};3)$
$(1.888_{-0.175}^{+0.081})\cdot 10^{-4}$
$m_{1}=\overline{m}_{s}(2\,\text{GeV})=0.00935\,\text{GeV}$
$m_{2}=\overline{m}_{b}(\overline{m}_{b})=4.183\,\text{GeV}$
$\chi_{T}(0;3)$ $(4.079_{-0.593}^{+0.121})\cdot 10^{-4}$
$\chi_{AT}(0;3)$ $(4.042_{-0.585}^{+0.118})\cdot 10^{-4}$
$\chi_{T}(Q_{b}^{2};3)$ $(1.893_{-0.175}^{+0.083})\cdot 10^{-4}$
$\chi_{AT}(Q_{b}^{2};3)$ $(1.884_{-0.174}^{+0.080})\cdot 10^{-4}$
$m_{1}=\overline{m}_{c}(\overline{m}_{c})=1.273\,\text{GeV}$
$m_{2}=\overline{m}_{b}(\overline{m}_{b})=4.183\,\text{GeV}$
$\chi_{T}(0;3)$ $(4.857_{-0.723}^{+0.173})\cdot 10^{-4}$
$\chi_{AT}(0;3)$ $(2.400_{-0.335}^{+0.078})\cdot 10^{-4}$
$\chi_{T}(Q_{b}^{2};3)$ $(2.248_{-0.215}^{+0.114})\cdot 10^{-4}$
$\chi_{AT}(Q_{b}^{2};3)$ $(1.404_{-0.149}^{+0.054})\cdot 10^{-4}$
Table 1: Central values (with all scales fixed to $\mu_{b}\equiv\overline{m}_{b}(\overline{m}_{b})=4.183\,\text{GeV}$) and corresponding uncertainties of the function $\chi_{T}(Q^{2};k)$ for $Q^{2}=\{0,Q_{b}^{2}\equiv\overline{m}_{b}^{2}(\overline{m}_{b})\}$ and different quark masses. See the text for further details.
We choose as reference scale $\mu_{b}\equiv\overline{m}_{b}(\overline{m}_{b})=4.183\,\text{GeV}$. The theoretical uncertainties are estimated by independently varying the four scales that enter our calculation: two associated with the quark masses, one with the strong coupling, and one with the current coupling. For each scale, the correlators are evaluated at three points, $\{\mu_{b}/2,\mu_{b},2\mu_{b}\}$, discarding extreme configurations in which two scales differ by more than a factor of two. This restricted independent variation results in $31$ distinct scale combinations. This prescription is a straightforward extension of the usual 7-point scale variation adopted for the case of two arbitrary scales to the case of four scales. There are several examples in the literature (see e.g. Refs. [20, 21, 22, 23, 24, 25]) where a similar extension to three scales was employed and found to yield good estimates of the scale uncertainties. The correlator currents are then consistently evolved to the reference scale $\mu_{b}$ using the renormalization group equation
$\frac{d\Pi_{T}}{d\ln{\mu^{2}}}=2\,\gamma_{T}\,\Pi_{T}\,$, (3.1)
where $\gamma_{T}$ is the anomalous dimension of the tensor current [26]:
$\frac{d(\bar{\psi}_{1}\sigma^{\mu\alpha}q_{\alpha}\psi_{2})}{d\ln{\mu^{2}}}$ $=\gamma_{T}\,\bar{\psi}_{1}\sigma^{\mu\alpha}q_{\alpha}\psi_{2}\,$, $\gamma_{T}$ $=\sum_{i=0}^{\infty}\gamma_{T,i}\bigg{(}\frac{\alpha_{s}}{\pi}\bigg{)}^{\!\!i+1},$ (3.2)
$\gamma_{T,0}$ $=-\frac{1}{3}\,$, $\gamma_{T,1}$ $=-\frac{181}{72}+n_{f}\frac{13}{108}\,.$
The solution of Eq. (3.1) is given by
$\Pi_{T}(q^{2};\mu_{b})$ $=\Pi_{T}(q^{2};\mu_{0})\exp\Bigg{\{}\int\limits_{\alpha_{s}(\mu_{0})}^{\alpha_{s}(\mu_{b})}\frac{2\gamma_{T}(\alpha)}{\beta(\alpha)}\,d\alpha\Bigg{\}}\,$, (3.3)
where [27, 28]
$\frac{d\alpha_{s}}{d\ln{\mu^{2}}}$ $=\>\beta\;,$ $\beta$ $=-\alpha_{s}\sum_{i=0}^{\infty}\beta_{i}\bigg{(}\frac{\alpha_{s}}{4\pi}\bigg{)}^{i+1},$ $\beta_{0}$ $=11-\frac{2}{3}n_{f},$ $\beta_{1}$ $=102-\frac{38}{3}n_{f},$ (3.4)
is the QCD beta function. This procedure ensures that the dominant scale dependence, already captured at NLO, largely cancels upon running, while the residual dependence provides a reliable estimate of the missing higher-order corrections.
We remark that the uncertainties on $\chi_{T}(Q^{2};k)$ are sizeable, with the lower uncertainty reaching up to $\sim 15\%$ and the upper uncertainty up to $\sim 5\%$. Nevertheless, these uncertainties are usually neglected in the context of unitarity bounds. A straightforward way to incorporate them is to adopt the upper values of $\chi_{T}(Q^{2};k)$, which constitutes the most conservative choice (see, e.g., Refs. [8, 29]). The magnitude of these uncertainties further motivates the necessity of an NNLO calculation of $\chi_{T}(Q^{2};k)$.
Figure 2: Comparison of our LO (green) and NLO (blue) results with lattice QCD data[7] (magenta) for different values of $m_{1}$.
[…]
In Fig. 2 we show our LO and NLO results for $\chi_{T}(0;3)$ and $\chi_{AT}(0;3)$. As seen in the figure, the inclusion of NLO corrections significantly reduces the uncertainties compared to LO. For the charm-quark case, we also compare with lattice QCD results of Ref. [7], which are consistent within the uncertainties. Note that our perturbative calculation and the lattice QCD results can be regarded as two independent calculations, and hence their results can be combined to achieve smaller uncertainties:
$\chi_{T}(0;3)$ $=(4.92\pm 0.22)\cdot 10^{-4}\,\text{GeV}^{-2}\,$, $\chi_{AT}(0;3)$ $=(2.39\pm 0.14)\cdot 10^{-4}\,\text{GeV}^{-2}\,.$ (3.5)
To obtain these averages, the central values of our results were shifted to the means of their asymmetric distributions, thereby rendering the uncertainties symmetric.
4 Summary and conclusion
We have presented the first complete analytical calculation of the vacuum-to-vacuum correlator of two quark tensor currents with two massive quarks, retaining the full momentum dependence. Our results extend previous studies by providing NLO corrections for arbitrary $q^{2}$, thereby filling a gap in the literature and correcting an inconsistency in the results of Ref. [6], which seemingly omitted the renormalization of the tensor current.
We provide all analytical expressions in machine-readable Mathematica format, facilitating their use in applications such as unitarity bounds and QCD sum rules. Using these expressions, we have presented numerical results for selected quark-mass values and demonstrated that the NLO corrections reduce the theoretical uncertainties compared to the LO predictions. The uncertainties, though reduced, remain non-negligible and should be accounted for in precision applications.
Our results show good agreement with a recent lattice QCD calculation [7], providing an independent cross-check and enabling a combination of perturbative and non-perturbative determinations to further improve precision. Finally, the magnitude of the remaining uncertainties motivates the calculation of NNLO corrections, which would be the natural next step toward fully controlled predictions for tensor-current correlators.
Overall, the work presented here provides essential theoretical input for dispersive analyses in flavour physics and represents a significant step forward in the precise determination of hadronic matrix elements involving tensor currents.
Acknowledgements
We thank the Cambridge Pheno Working Group and for insightful discussions. NG is grateful to D. van Dyk for valuable input on the ideas underlying this work. NG is thanks the authors of Ref. [6] for providing their results in a machine readable form and for helpful discussion.
This work has been partially supported by STFC consolidated grants ST/T000694/1 and ST/X000664/1. EL receives further support from the Alfred Kordelin Foundation.